The average length of time for students to enroll has been 50 minutes. If a random sample of 12 students incurred an average enrolment time of 42 minutes with a standard deviation of 11.9 minutes under a new enrolment system, test the hypothesis that the population mean is now less than 50 using a 0.05 level of significance. Assume that the length of time of the population is normal.


1
Expert's answer
2022-03-15T17:05:24-0400

Solution.

The null hypothesis H0: "\mu=50"

The alternative hypothesis H1: "\mu<50"


Sample size "n=12"

Sample mean "\bar X =42"

Sample standard deviation "\sigma=11.9"

Significance level "\alpha =0.05"


Since the alternative hypothesis is that "\mu" is Less than this is a one tailed hypothesis test.


Since the sample size is less than 30 and the population standard deviation is not known, we use the T test


"df =n-1=12-1=11"


Since "\alpha = 0.05" , T "=-1.7959"


"t_c =\dfrac{\bar X -\mu_0 }{S/\sqrt n}=\dfrac {42-50}{11.9/\sqrt12 }"


"t_c =-2.3288"


Since "t_c<T" we reject the hypothesis.


Since the null hypothesis is rejected, there is sufficient evidence to conclude that the population mean is now less than 50 at 0.05 level of significance.